# Quantum memory at nonzero temperature in a thermodynamically trivial system

### v1

@@ 4 Measurement-free quantum error correction
Due to the spatial non-locality of the $\mathsf{k}$-local LDPC codes, it may be challenging to simply “let them thermalize” with a cold environment to build a quantum memory. Nevertheless, let us now explain a very useful and practical alternative, based on “measurement-free quantum error correction” (MFQEC) Ahn et al. (2002); Sarovar and Milburn (2005), in which the user can only apply unitary gates and a specific “reset” dissipative quantum channel to the physical qubits. More formally, given density matrix $\rho$ for the physical qubits, we can only apply $\rho\rightarrow U\rho\,U^{\dagger}$ for unitary $U$, or apply $\rho\rightarrow R_{i}[\rho]$ where channel $R_{i}$ acts on single qubit $i$ as $R_{i}[\rho_{i}]=|0\rangle\langle 0|$; here we defined $Z|0\rangle=|0\rangle$. In addition to these processes, we also have errors occurring in the system, which we take to be independent and few-qubit, albeit occurring at a constant rate per qubit. The strategy behind MFQEC is to introduce ancilla qubits in addition to the $N$ physical qubits of the code. At the beginning of each error correction round, we apply the reset channel such that the state of the system is $|\psi_{\mathrm{total}}\rangle=|\psi_{\mathrm{phys}}\rangle\otimes|\mathbf{0}\rangle_{\mathrm{anc}}$. The role of the ancilla qubits is best understood as storing the measurement outcomes that would have occurred in conventional quantum error correction, which proceeds by measuring all the parity-checks (syndromes) and deducing where errors occurred based on these outcomes. In MFQEC, rather than measuring a stabilizer, we instead apply CNOT gates to encode the parity of the stabilizer into the state of the ancilla qubit (so long as CNOT is applied error-free). Then, we apply gates to the code conditioned on the state of the ancilla – i.e., we error correct based on syndromes. Using Stinespring’s formulation of measurement Stinespring (1955); Friedman et al. (2022) wherein measurement outcomes are stored in auxiliary qubits in Hilbert space such that the measurement process becomes unitary, a round of MFQEC is mathematically equivalent to a round of conventional QEC.
Still, from an experimental perspective, MFQEC can be highly desirable. Once we have a memoryless local decoder, implementing it via MFQEC will only require a constant number of ancilla bits per physical bit to store the measurement outcomes necessary for error correction. For one thing, once the error correction has been applied, we no longer need memory of those measurement outcomes, and so the ancilla qubits can be reset. More importantly, and in contrast to single-shot decoding strategies that use global syndrome information, we only need a finite-depth circuit to implement the passive feedback via MFQEC.
Let us illustrate how to perform MFQEC for our Gibbs sampler, using $\leq 3$-qubit gates, and reset. For simplicity, we focus only on the step that “corrects” $X$ errors on qubit $i$, as the story is analogous for $Z$ errors and other qubits. Let ancilla qubits $a_{1},\ldots,a_{\Delta_{C}}$ store the “measurement outcomes” for the $\Delta_{C}$ $Z$-checks $C_{1},\dots,C_{q}$ containing physical qubit $i$. The first step of MFQEC is to apply the unitary
where $\mathsf{CNOT}_{i\rightarrow j}=\frac{1+Z_{i}}{2}+\frac{1-Z_{i}}{2}X_{j}$. In the second step, we apply the $\mathsf{SORT}_{\Delta_{C}}$ channel, which shuffles the ancilla qubits such that the zeros are to the left and the ones are to the right. We provide an implementation for $\mathsf{SORT}$ in Appendix D using $\leq 3$-qubit gates and reset. In order to realize a perfect Gibbs sampler, we apply a coherent gate that implements the following controlled-unitary using one additional ancilla qubit $c$, which is again initialized in state $|0\rangle$:
where we choose each angle $\theta_{j}$ such that
As a practical matter, we expect that our Gibbs sampler can be modified to never introduce unwanted errors, and that the lifetime of logical information should only increase. This also simplifies the circuit above (see Appendix D). More importantly, we note that our protocol is fault-tolerant: it works even if there is noise during the MFQEC process itself. As an explicit example of this, we show in Appendix D that if the overall single-qubit error rates are sufficiently small (but finite), it is possible to Gibbs sample in the presence of errors during the MFQEC circuit.
@@ 5 Outlook
We have proved that random classical and quantum LDPC codes are constant-rate quantum memories, in which $N$ physical (qu)bits which undergo local, memoryless, time-reversal-symmetric Gibbs-sampling dynamics at low temperature, thus protecting $\mathrm{O}(N)$ logical (qu)bits for infinite time in the thermodynamic ($N\rightarrow\infty$) limit. This property holds even as the codes are thermodynamically trivial – they have no thermodynamic phase transitions at nonzero temperature. Our result provides a deeper understanding of the surprising connections between statistical physics and error correction: contrary to prior intuition, thermal phase transitions are unnecessary for passive error correction, which can be performed by sampling the thermal Gibbs state, in either classical or quantum codes.
@@
The self-correcting capability of LDPC codes has a number of intriguing implications and connections with other research thrusts in physics. The LDPC codes we described in this paper are arguably the simplest (albeit high-dimensional) examples of models exhibiting dynamical state/Hilbert space shattering in the thermodynamic limit: the system is dynamically trapped in an exponentially small region of the state space, even in the absence of any protecting symmetry or even simply frustration or Hilbert space constraints. Our result thus makes a sharp connection between error-correcting codes and recent models of ergodicity breaking in statistical physics Sala et al. (2020); Khemani et al. (2020); Hart and Nandkishore (2022); Stephen et al. (2024); Stahl et al. (2023); Han et al. (2024).
@@
Since no locally-testable good quantum LDPC codes ($c^{3}$-qLTCs) are known yet, we conjecture that the non-extensive redundancy of our quantum hypergraph product codes is also shared by the existing good (constant-rate $K\sim N$ and linear-distance $D\sim N$) quantum LDPC codes Panteleev and Kalachev (2022); Leverrier and Zémor (2022); Dinur et al. (2022b). If true, our proof immediately generalizes to such models, and implies that circuit complexity of preparing states is not a universal characteristic of a thermodynamic phase – namely, that the circuit depth needed to prepare a state at various energy densities will transition from finite to infinite (in the thermodynamic limit) without crossing any thermal phase transition. After all, such codes have the no-low-energy-trivial-states (NLTS) property Freedman and Hastings (2013); Anshu et al. (2023) – arbitrary eigenstates of a Hamiltonian, up to a finite energy density, are topologically ordered: they cannot be constructed from a product state via a finite-depth circuit. Nonetheless, extensive redundancy is required for phase transitions at nonzero temperature! This conclusion is rather surprising: we currently define two pure states to be in the same phase only when they are related by a finite-depth unitary circuit Chen et al. (2010). The situation for mixed states is less clear, see Sang et al. (2023) for recent developments. Evidently, NLTS would thus reveal two mixed states in the same thermodynamic phase which simultaneously violate the above criterion. A better understanding of such a discrepancy is an interesting problem for the future.
Lastly, let us remark on the relevance of our results for neutral atom quantum computing Saffman et al. (2010); Kaufman and Ni (2021); Cong et al. (2022); Bluvstein et al. (2022); Jenkins et al. (2022); Bluvstein et al. (2023). (1) In this platform, the reset operation can be about 100 times faster than mid-circuit measurement and feedback Lis et al. (2023); Norcia et al. (2023); Huie et al. (2023). The circuit depth needed for MFQEC is a constant-factor $\lesssim 3$ larger than that needed for QEC based on mid-circuit measurement and feedback (at the cost of adding $\mathrm{O}(N)$ extra ancillas), so we expect that a passive decoder would operate faster. The primary disadvantage of the passive decoder – a possibly slow implementation of nonlocal atomic motion – would be just as problematic for the syndrome measurements themselves. (2) HGP codes are well-suited to the hardware capabilities of atom arrays Xu et al. (2023); Hong et al. (2023), in which acousto-optical deflectors can perform row and column permutations nonlocally in space relatively quickly (at least, for near-term system sizes). Hence, spatial nonlocality of the LDPC code is not necessarily prohibitive of its implementation in the absence of a large quantum-repeater network Fowler et al. (2010); Azuma et al. (2023). (3) Neutral atom platforms have increased their number of physical qubits Bluvstein et al. (2023) more rapidly than other platforms, which suggests that the additional ancilla qubit overhead for the fastest passive decoding strategy is less burdensome. Lastly, our numerical simulations suggests that the rigorous bounds on threshold error rates are far more stringent than needed in practice. While a detailed cost-benefit analysis of measurement-free vs. syndrome-based error correction in this platform remains to be performed, and the precise implementation of any optimal (possibly hybrid passive/active) error-correcting protocol remains to be found, we are optimistic that passive decoding of LDPC-based quantum memories is particularly suitable for neutral atom platforms and can play an important role in realizing fault-tolerant quantum computers in the future.
@@ Acknowledgements
We thank Oliver Hart, Mike Hermele, Vedika Khemani, Anthony Leverrier, Marvin Qi, Rahul Nandkishore, Zohar Nussinov and Charles Stahl for helpful discussions, Xun Gao for a careful reading of the manuscript, and especially Adam Kaufman for many insights on neutral atoms. This work was supported by the Alfred P. Sloan Foundation through Grant FG-2020-13795 (AL), the Air Force Office of Scientific Research via Grant FA9550-21-1-0195 (YH, JG, AL), and the Office of Naval Research via Grant N00014-23-1-2533 (YH, AL).
@@ A.1 Linear parity-check codes
We say that $H$ is $(\Delta_{B},\Delta_{C})$ LDPC (low-density parity-check) if the maximum Hamming weights of its rows and columns are $\Delta_{C},\Delta_{B}=O(1)$ respectively. In other words, every check acts on a constant number of physical bits, and every bit only participates in a constant number of checks. If both $k=\Theta(n)$ (constant rate) and $d=\Theta(n)$ (linear distance), then we say the code is “good”. For most of this paper, our interest is on LDPC codes, although some of the results we derive below are not specific to this case.
@@
We can define a $(\Delta_{B},\Delta_{C})$-LDPC expander code $\mathcal{C}_{\mathcal{G}}$ by identifying the left-nodes $B$ with physical bits and the right-nodes $C$ with parity checks. The parity-check matrix $H$ is given by the bipartite adjacency matrix of $\mathcal{G}$. Using the above definition of a bipartite expander, one can immediately derive lower bounds on the code parameters of $\mathcal{C}$.
@@
Now that we have reviewed how bipartite expander graphs can lead to good LDPC codes, the problem boils down to constructing bipartite graphs with the required expansion properties. Explicit constructions with $\delta<1/2$ are notoriously difficult; see, however, Capalbo et al. (2002). Nonetheless, if one randomly constructs a bipartite graph using probabilistic methods, it will often exhibit the required expansion with high probability. We focus on regular ensembles, where the target degree distributions are singular around $\Delta_{B},\Delta_{C}$; we call such an ensemble the $(\Delta_{B},\Delta_{C})$-LDPC ensemble. The success of the random construction can be quantified by the following result.
@@
An interesting property of non-redundant codes is that all $2^{m}=2^{n-k}$ syndromes have a corresponding physical error up to the action of a codeword. To see this fact, first note that $H$ has full rank because all its rows are linearly independent. Because row and column ranks are equal, we can thus select $m=n-k$ linearly independent columns of $H$ to form a new, invertible $m\times m$ matrix $\tilde{H}$. For any syndrome $\mathbf{s}$, a corresponding error $\mathbf{e}(\mathbf{s})$ can be computed using $\mathbf{e}(\mathbf{s})=\tilde{H}^{-1}\mathbf{s}$. As a consequence, there exists error vectors which correspond to a single flipped parity check (excitation). Nonetheless, from the expansion properties of the underlying Tanner graph, we can show that such low-energy excitations necessarily have extended support, as formalized in the following Corollary.

### v2

@@ 4 Measurement-free quantum error correction
Due to the spatial nonlocality of the $\mathsf{k}$-local LDPC codes, it may be challenging to simply “let them thermalize” with a cold bath: physical qubits must be arranged in three-dimensional space, and the most efficient way to implement the LDPC interactions may require dynamical motion of the physical qubits. Nevertheless, the existence of passive decoders enable new experimental paradigms for error correction when compared to spatially local (e.g. surface) codes. One strategy uses “measurement-free quantum error correction” (MFQEC) Ahn et al. (2002); Sarovar and Milburn (2005), in which the user can only apply unitary gates and a specific “reset” dissipative quantum channel to the physical qubits. In the many-body setting, MFQEC typically proceeds by engineering local dissipation Pastawski et al. (2011). More formally, given density matrix $\rho$ for the physical qubits, we can only apply $\rho\rightarrow U\rho\,U^{\dagger}$ for unitary $U$, or apply $\rho\rightarrow R_{i}[\rho]$ where channel $R_{i}$ acts on single qubit $i$ as $R_{i}[\rho_{i}]=|0\rangle\langle 0|$; here we defined $Z|0\rangle=|0\rangle$. In addition to these processes, we also have errors occurring in the system, which we take to be independent and few-qubit, albeit occurring at a constant rate per qubit. The strategy behind MFQEC is to introduce ancilla qubits in addition to the $N$ physical qubits of the code. At the beginning of each error correction round, we apply the reset channel such that the state of the system is $|\psi_{\mathrm{total}}\rangle=|\psi_{\mathrm{phys}}\rangle\otimes|\mathbf{0}\rangle_{\mathrm{anc}}$. Next, we encode the state of stabilizers into the $Z$-eigenstate ancilla qubits. In conventional quantum error correction, we would then measure these ancilla qubits in the $Z$-basis, and apply feedback based on the outcomes. In MFQEC, we instead just directly apply feedback to the physical qubits based on multi-qubit gates which couple them again to the ancillas. After this feedback is applied, we reset the ancillas. Using Stinespring’s formulation of measurement Stinespring (1955); Friedman et al. (2022) wherein measurement outcomes are stored in auxiliary qubits in Hilbert space such that the measurement process becomes unitary, a round of MFQEC is mathematically equivalent to a round of conventional QEC.
If passive decoders rely only on local feedback, then the decoding circuit for MFQEC has finite depth and requires only finite ancillas per physical qubit to implement. In Appendix E, we discuss details of how to Gibbs sample with MFQEC and show that such a Gibbs sampler implies the existence of a fault-tolerant decoder. The crucial observation is that a Gibbs sampler is local and always introduces errors into the system with non-zero probability. If the net error rate (including errors in syndrome measurement and decoding) $p_{\mathrm{error}}\leq\mathrm{e}^{-2\beta\Delta_{C}}$, where $\beta>\beta_{\mathrm{c}}$ implies that we are below the critical temperature where Gibbs sampling is slow, and $\Delta_{C}$ is the maximal number of stabilizers that a single qubit error can flip, then we can always add additional errors into the system (for qubits where some stabilizers were already violated) such that our error correction scheme realizes a perfect Gibbs sampler at low temperature, which we have proven is a passive decoder.
In practice, we expect the Gibbs sampler can be replaced by a greedy decoder, such as bit-flip Sipser and Spielman (1996), which always tries to lower the energy of the system. Such a local decoder will not in general drive the system exactly to its logical codeword. When information is to be retrieved, it can be done using conventional decoders. Gibbs sampling thus serves as a particular kind of “single-shot decoder”, where a single round of noisy syndrome measurement suffices to suppress errors to a correctable level Bombin (2015); Campbell (2019). Single-shot decoders exist for a variety of quantum LDPC code families Quintavalle et al. (2021); Higgott and Breuckmann (2023), with some decoders also requiring only local syndrome information Fawzi et al. (2018); Gu et al. (2023). The single-shot decoders in Quintavalle et al. (2021); Higgott and Breuckmann (2023) still require global syndrome information, and so direct implementation with MFQEC is not practical. $\mathsf{k}$-locality in passive decoding is crucial to make MFQEC tractable; slow, computationally expensive decoding can be done once information is retrieved out of memory. Gibbs samplers lead to MFQEC decoders with significantly lower circuit depth than the single-shot decoders in Fawzi et al. (2018); Gu et al. (2023).
Perhaps the platform where MFQEC-based passive decoding is most desirable is neutral-atom quantum computing Saffman et al. (2010); Kaufman and Ni (2021); Cong et al. (2022); Bluvstein et al. (2022); Jenkins et al. (2022); Bluvstein et al. (2023). (1) In this platform, the reset operation can be about 100 times faster than mid-circuit measurement and feedback Lis et al. (2023); Norcia et al. (2023); Huie et al. (2023). The circuit depth needed for MFQEC is a constant-factor $\lesssim 3$ larger than that needed for QEC based on mid-circuit measurement and feedback (at the cost of adding $\mathrm{O}(N)$ extra ancillas), so we expect that a passive decoder would operate faster. The primary disadvantage of the passive decoder – a possibly slow implementation of nonlocal atomic motion – would be just as problematic for the syndrome measurements themselves. (2) HGP codes are well-suited to the hardware capabilities of atom arrays Xu et al. (2023); Hong et al. (2023), in which acousto-optical deflectors can perform row and column permutations nonlocally in space relatively quickly (at least for near-term system sizes). Hence, spatial nonlocality of the LDPC code is not necessarily prohibitive of its implementation in the absence of a large quantum-repeater network Fowler et al. (2010); Azuma et al. (2023). (3) Neutral atom platforms have increased their number of physical qubits Bluvstein et al. (2023) more rapidly than other platforms, which suggests that the additional ancilla qubit overhead for the fastest passive decoding strategy is less burdensome. Lastly, our numerical simulations suggest that the rigorous bounds on threshold error rates are far more stringent than needed in practice. A detailed cost-benefit analysis of measurement-free vs. syndrome-based error correction in this platform remains to be performed.
@@ 5 Outlook
We have proved that classical and quantum LDPC codes with linear confinement are self-correcting memories, in which $N$ physical (qu)bits which undergo local, memoryless, time-reversal-symmetric Gibbs-sampling dynamics at low but nonzero temperature, protecting $\mathrm{O}(N)$ logical (qu)bits for infinite time in the thermodynamic ($N\rightarrow\infty$) limit. This property holds even as the codes are thermodynamically trivial – they have no thermodynamic phase transitions at nonzero temperature. Our result provides a deeper understanding of the surprising connections between statistical physics and error correction: contrary to prior intuition, thermal phase transitions are unnecessary for passive error correction, which can be performed by sampling the thermal Gibbs state, in either classical or quantum codes. Although our analysis is of random codes, see Capalbo et al. (2002); Golowich (2024) for explicit constructions with linear confinement. We do not know whether these explicit codes also share the same thermodynamic triviality as their random counterparts.
An immediate application of our results is a local Markovian (memoryless) decoder based on Gibbs sampling that can passively protect quantum information in (expander) hypergraph product codes, which presents an alternative “decoding” strategy for error correction compared to existing ones Leverrier et al. (2015); Panteleev and Kalachev (2021); Quintavalle and Campbell (2022); Delfosse et al. (2022); note that while our Gibbs sampler is not guaranteed to return the noisy state exactly back to the codespace, it is guaranteed to keep the noisy state sufficiently close to its initial codeword so that eventually one can successfully decode the final state using one of the existing (full) decoders. Especially for MFQEC, the complexity of the thermal decoder is substantially reduced relative to prior examples.
@@
The self-correcting capability of LDPC codes has a number of intriguing implications and connections with other research thrusts in physics. The LDPC codes we described in this paper are arguably the simplest (albeit high-dimensional) examples of models exhibiting dynamical state/Hilbert space shattering in the thermodynamic limit: the system is dynamically trapped in an exponentially small region of the state space, even in the absence of any protecting symmetry or even simply frustration or Hilbert space constraints. Our result thus makes a sharp connection between quantum error-correcting codes and recent models of ergodicity breaking in (quantum) statistical physics Sala et al. (2020); Khemani et al. (2020); Hart and Nandkishore (2022); Stephen et al. (2024); Stahl et al. (2023); Han et al. (2024). In the above models, the systems are closed and evolve according to their own Hamiltonians. The dynamics we study, on the other hand, are local spin flips resembling open system dynamics. For closed-system dynamics of LDPC codes, we refer the reader to a related model exhibiting explicit eigenstate localization Yin et al. (2024), which is an even more drastic breakdown of ergodicity (to infinite time). The classical LDPC code Hamiltonians we have described Montanari and Semerjian (2006a, b) are also qualitatively different from many common models of spin glasses Edwards and Anderson (1975); Sherrington and Kirkpatrick (1975); Mezard et al. (1987) such as the Sherrington-Kirkpatrick model, and random instances of many NP-hard optimization problems Krzakala et al. (2007), which have analogous ergodicity-broken dynamical regimes at low temperature, yet also have frustration and thermodynamically-detectable phase transitions.
@@
The existence of locally testable, good quantum LDPC codes ($c^{3}$-qLTCs) is still an open question. These codes are nonlocal generalizations of the 4D toric code, similar to $c^{3}$-LTCs and the 2D Ising model. Importantly, they will contain linear dependencies amongst their parity checks in a way such that all low-energy states will necessarily be close to codewords, with “close” being quantified by the codes’ soundness functions Aharonov and Eldar (2015). For example, in the 4D toric code, small error clusters form closed loops of violated checks whose area scales quadratically with its perimeter, which defines the soundness function. Thus, just like the 4D toric code, we expect $c^{3}$-qLTCs to possess genuine thermodynamic phase transitions. Historically, since all the prototypical examples of passive quantum memories contained thermodynamic phase transitions, the natural object to study for a constant-rate, passive quantum memory would be a qLTC. In contrast to this intuition, we have shown that existing (non-LTC) constructions of constant-rate quantum codes, namely the hypergraph product codes, possess the sufficient ingredients for passive error correction. In addition, since the existing good (constant-rate $K\sim N$ and linear-distance $D\sim N$) quantum LDPC codes Panteleev and Kalachev (2022); Leverrier and Zémor (2022); Dinur et al. (2022b) are likely not locally testable, we conjecture that the non-extensive redundancy of random hypergraph product codes is also shared by these code families. If true, our proof immediately generalizes to such models and implies that the circuit complexity of preparing states is not a universal characteristic of a thermodynamic phase – namely, that the circuit depth needed to prepare a state at various energy densities will transition from finite to divergent without crossing any thermal phase transition. After all, such codes exhibit the no-low-energy-trivial-states (NLTS) property Freedman and Hastings (2013); Anshu et al. (2023) – arbitrary low-energy states of a Hamiltonian, up to a finite energy density, are topologically ordered: they cannot be constructed from a product state via a finite-depth circuit. Based on a similar intuition as passive memories mentioned above, it used to be expected that NLTS was closely related to local testability of qLTCs Eldar and Harrow (2017); Panteleev and Kalachev (2022); the surprising result that local testability was not needed Anshu et al. (2023) may be deeply related to our own observation that thermodynamics and self-correction are not necessarily related. We currently define two pure states to be in the same phase only when they are related by a finite-depth unitary circuit Chen et al. (2010), with the analogy for mixed states less clear Rakovszky et al. (2024); Sang et al. (2023); Chen and Grover (2023). We conclude that NLTS in non-redundant codes would thus reveal two states in the same trivial thermodynamic phase, yet which cannot be efficiently prepared from each other. A better understanding of such a discrepancy is an interesting problem for the future, and may sharpen our understanding of how to define phases of matter in quantum many-body systems.
@@ Acknowledgements
We thank Oliver Hart, Mike Hermele, Vedika Khemani, Anthony Leverrier, Rahul Nandkishore, Zohar Nussinov, Marvin Qi and Charles Stahl for helpful discussions, Xun Gao for a careful reading of the manuscript, and especially Adam Kaufman for many insights on neutral atoms. This work was supported by the Alfred P. Sloan Foundation through Grant FG-2020-13795 (AL), the Air Force Office of Scientific Research via Grants FA9550-21-1-0195 and FA9550-24-1-0120 (YH, JG, AL), and the Office of Naval Research via Grant N00014-23-1-2533 (YH, AL).
@@ A.1 Linear parity-check codes
We say that $H$ is $(\Delta_{B},\Delta_{C})$ LDPC (low-density parity-check) if the maximum Hamming weights of its rows and columns are $\Delta_{C},\Delta_{B}=O(1)$ respectively. In other words, every check acts on a constant number of physical bits, and every bit only participates in a constant number of checks. If both $k=\Theta(n)$ (constant rate) and $d=\Theta(n)$ (linear distance), then we say the code has asymptotically “good” parameters. For most of this paper, our interest is on LDPC codes, although some of the results we derive below are not specific to this case.
@@
Def. A.1 is also referred to as (one-sided or two-sided) lossless expansion in the literature, with $\delta$ quantifying the amount of “loss” from maximum expansion. We can define a $(\Delta_{B},\Delta_{C})$-LDPC expander code $\mathcal{C}_{\mathcal{G}}$ by identifying the left-nodes $B$ with physical bits and the right-nodes $C$ with parity checks. The parity-check matrix $H$ is given by the bipartite adjacency matrix of $\mathcal{G}$. Using the above definition of a bipartite expander, one can immediately derive lower bounds on the code parameters of $\mathcal{C}$.
@@
Now that we have reviewed how bipartite expander graphs can lead to good LDPC codes, the problem boils down to constructing bipartite graphs with the required expansion properties. Explicit, algebraic constructions of lossless expanders are known to exist Capalbo et al. (2002); Golowich (2024) with the required properties. On the other hand, it also turns out that random bipartite graphs are lossless expanders with high probability Richardson and Urbanke (2008). For simplicity of analysis, we will focus on regular ensembles, where the target degree distributions are singular around $\Delta_{B},\Delta_{C}$; we call such an ensemble the $(\Delta_{B},\Delta_{C})$-LDPC ensemble. The success of the random construction can be quantified by the following result.
@@
An interesting property of non-redundant codes is that all $2^{m}=2^{n-k}$ syndromes have a corresponding physical error up to the action of a codeword. To see this fact, first note that $H$ has full rank because all its rows are linearly independent. Because row and column ranks are equal, we can thus select $m=n-k$ linearly independent columns of $H$ to form a new, invertible $m\times m$ matrix $\tilde{H}$. For any syndrome $\mathbf{s}$, a corresponding error $\mathbf{e}(\mathbf{s})$ can be computed using $\mathbf{e}(\mathbf{s})=\tilde{H}^{-1}\mathbf{s}$. As a consequence, there exists error vectors which correspond to single-flipped parity checks (excitations). Nonetheless, from the expansion properties of the underlying Tanner graph, we can show that such low-energy excitations necessarily have extended support. Intuitively, one can think of a low-energy excitation as a codeword of a modified code consisting of the original code but with the flipped check removed. Since removing a single check is unlikely to change the global expansion properties of the Tanner graph, the modified code will also have macroscopic distance, which implies that our low-energy excitation is far from all original codewords. We formalize this argument in the following Corollary.